OpenStudy (anonymous):

What is the solution to the equation? 2 - z/3 = 12

OpenStudy (anonymous):

A. z = –42 B. z = –34 C. z = –30 D. z = 42

OpenStudy (anonymous):

@alesia/bailey

OpenStudy (mathivh):

In steps: -z/3=12-2 (-2) -z/3=10 (simplify) z/3=-10 (*-1) z=-30 (*3)

OpenStudy (mathivh):

So the answer is C; (-30).

OpenStudy (anonymous):

Thank you